celticcat:

This one as really got me foxed! The book I'm looking at says you use another form of the FLT If x^(p-1) not = 1 mod p then p is not prime. Then it asks you to show that 66,013 is not prime. Such a big number and I find modular arithmetic pretty difficult!

celticcat:

I can understand how you could prove it for say 39. We would show that 2^38 NOT = 1 mod 39 and make the arithmetic easier by writing 2^38 = 2^36 . 2^2.

celticcat:

i guess you could start by writing 2^66012 = 6^65536 . 2^256.2^128.2^64.2^16.2^12 mod 66013 but you would need a big number crunching computer to work that out!!

celticcat:

the last 5 expressions could be handled by a good on-line calculator I guess.

will:

im sorry i cant get the answer it keeps going in to a error breakdown'

celticcat:

and you would use the relation rs mod p = r mod p * s mod p

celticcat:

you using a calculator?

will:

yes'

celticcat:

you can get a value for 2^32 mod 66013 using a calculator which is 497 the to get 2^64 you square 497 the repeat to process of finding the remainder

celticcat:

2^65536 is the problem

celticcat:

you might be able to simplify that using mod arithmetic Ill have to hit the books, i guess

celticcat:

wow!!!

Ultrilliam:

That is this: 2^65536

celticcat:

how did you get that?

Ultrilliam:

This server has 1152 cores, I borrowed 1 to do the calculation lol

Ultrilliam:

Something tells me you probably didn't want that massive number though xD

celticcat:

now we need to divide that by 66013 and find the remainder

celticcat:

lol no!

Ultrilliam:

Oh boy, well I gotta run to the bank, can you wait a moment? LOL

Ultrilliam:

By this server I meant the QC server xD

celticcat:

there must be an easier way to do this!!!!

Ultrilliam:

Probably xD

celticcat:

Hero messaged me asking what the textbook was, Maybe he has some ideas

celticcat:

talk to you later

sillybilly123:

reckon there is a brute-force way to do it which doesn't crash the calculator You can factor: \(66012 = 2^2 * 3 * 5501\) \(\implies 2^{66012} = 2^{4 * 3 * 5501} = \left (2^{ 5501}\right)^{12}\) Converting to binary: \(5501 \equiv {1010101111101}_2\) so \(5501 = 2^0 + 2^2 + 2^ 3 + 2^ 4 + 2^ 5 + 2^ 6 + 2^8 + 2^{10} + 2^{12}\) IE: \( 2^{5501} = 2^{2^0 + 2^2 + 2^ 3 + 2^ 4 + 2^ 5 + 2^ 6 + 2^8 + 2^{10} + 2^{12}}\) \( = 2 * 2^4 * 2^8 * 2^{ 16} * 2^{32} * 2^{ 64} * 2^{ 256} * 2^{ 1024} * 2^{ 4096}\) The mod of the product is the product of the mods, simple ones first: \(2 \mod 66013 = 2 \) \(2 ^4 \mod 66013 = 16 \) \(2 ^8 \mod 66013 = 256 \) if you work through all of that meticulously, remembering that \(a^k = \alpha ^ k \mod m\) and \(ab = \alpha \beta \mod m\), i think you can grind out a result. the site is already slooooooow so those previous posts do not help

Ultrilliam:

"slooooooow" But the site is fast ._.

Ultrilliam:

if it's slow it's client side ._.

Ultrilliam:

Oh yea, latex does take a while to start loading... I really need to work on ways to fix that, such as pre-loading the extensions and such...

Ultrilliam:

I thought you were talking about the site, not LaTeX

Ultrilliam:

There, after some work, mathjax now idly pre-loads assets for mathml upon page load, and instead of queuing latex to be loaded, i made it forcefully load the content, so as a result, LaTeX should now load faster

Ultrilliam:

still takes about 5-10 seconds to finish loading mathml in the background, I'll need to figure out why and fix that at some point however

Ultrilliam:

there we go :P

Ultrilliam:

Fast enough for you now? @sillybilly123

Ultrilliam:

(you might need to clear your cache)

sillybilly123:

thx U, i think you might be right 😊 celtic: there is an easy way to factor 66013. Fermat's factorization method http://openstudy.com/updates/56ae016be4b0da09eb82f8c0 in fact that was asking the exact same question i think.

celticcat:

yea Tanks billy. However they wanted you to use FLT. i finally used brute force and showed that it was not = 1 mod 66013. I found 2^65536 by repeating the calculation 2^256 --> 2^512 up to 2^65536

celticcat:

i had to use the WEb 2 online calculator which can handle very large numbers

celticcat:

Yeah Fermat's Factorization is much easier!